# Repeat After Me: Transformers are Better than State Space Models at Copying Transformers are Better than State Space Models at Copying

### v1

@@ 2.2 Transformers can copy inputs of exponential length
In this section we show that transformers can implement the copy operation for input sequences with length exponential in the number of heads. Namely, we construct a transformer with two blocks that gets small error on the copy task.
@@ 3.1 Experimental setup
We now provide a brief overview of our experimental setup. Further details may be found in Appendix A.
@@ 3.2 Data efficiency on the copy task
Results for this experiment are shown in 0(a). Clearly, there is a large gap between the transformers and GSSMs. We observe that the transformers need 100x less samples than the best GSSMs to learn the copy task.
Note that the sharp changes in accuracy displayed in 0(a) are due to the log-scaled x-axis and choice of string-level accuracy as a metric. In 8(a), we report the character-level accuracy, which yields smoother curves demonstrating the learning process of GSSMs. Regarding LSTMs, we find that they do not manage to learn on length-300 strings even at the character level. In 8(b), we show that LSTMs are able to learn to copy on shorter strings and that string length is the bottleneck.
@@ 3.3 Length generalization on the copy task
Here, we train all models on sequences of $\leq 50$ tokens, and test them on sequences of up to $1000$ tokens, reporting string-level accuracy. As seen in 0(b), all models are able to (eventually) solve the task in-distribution on lengths of $\leq 50$, but transformer-based models display much better generalization to longer inputs compared to GSSMs. Namely, we observe that the performance of the GSSMs (LSTM and MAMBA) drops to zero almost immediately when increasing the input length, while the performance of transformers decays much more gradually with length.
@@
When looking at the relative performance of different transformer models in 0(b), it becomes clear that the positional encoding is important to length generalization. Specifically, the ALiBi and NoPE transformers dramatically outperform the RoPE model on longer inputs. This is likely because the sinusoidal embeddings of RoPE create a more dramatic change than the decay of ALiBi or NoPE when we go to longer inputs.
@@
To test our understanding of how transformers learn to copy, we now consider swapping in the Hard-ALiBi positional encoding that we used in our theoretical construction of hash-based copying (introduces in Subsection 2.2 and illustrated in Figure 8 in the Appendix). 0(b) shows that a transformer trained with Hard-ALiBi embedding on sequences of length $\leq 50$ achieves almost perfect length generalization up to sequences of length 1000. Note that this is well beyond the context length ever encountered in training.
n-gram length:  2    3    4    5    6     7    8
@@ 3.4 Transformers learn to use n-gram hashing
Next, we attempt to determine whether the transformer trained on the copy task indeed applies the mechanism of storage and retrieval of n-grams. To do this, we evaluate the performance of a transformer with Hard-ALiBi positional encoding trained on the copy task when tested on a distribution of examples that intentionally contains duplicate n-grams. That is, we draw uniform sequences of tokens, and then randomly replace some n-gram with another n-gram that already appears in the sequence, such that each example always contains two copies of the same n-gram (typically followed by a different token). We use the Hard-Alibi model here since it performs the best for the copy task as showed in 0(a). Figure 4 shows the performance of the transformer for different choices of $n$. We observe that the transformer maintains roughly the same accuracy for $n\leq 4$, but that its accuracy starts dropping when the inputs contains duplicate sequences of 5 or more tokens. This suggests that the transformer relies on something like 5-gram retrieval to do the copy task.
@@ 4.1 Setup
We compare these models by measuring their performance while varying the input instance length and consider two types of tasks: copy-based and information retrieval tasks. The copy-based tasks consist of presenting a random text to the model and asking it to copy the text. In the information retrieval tasks, we provide a text to the model and ask it a related question. These retrieval tasks can be seen as ”selective copy”, since the model needs to copy a small chunk of the input text in order to respond to the question. To measure performance, we use the string-level accuracy in all the experiments except in 6(c) where we consider question answering and thus report the F1 score. We evaluate the models over 10 batches of size 64 for all the tasks except for question answering where we evaluate over 50 questions because the number of questions with a given context length is limited. Further details are in Appendix A.
@@ 4.2 Copying the input context
We first observe that pre-trained transformers outperform pre-trained GSSMs at copying long natural language strings. In 6(a), we randomly sample strings from the C4 dataset (Raffel et al., 2020) with varying number of tokens. Our prompt consists of two copies of the sampled string plus the first word of the string and we expect the model to complete the third copy. Even the smallest transformer model dramatically outperforms the largest GSSM. This happens even though the large GSSMs have enough bits in the state variable to potentially store the context. This confirms the idea that this is an architectual bias of transformers that makes it easier for them to copy from the context.
Unlike strings of tokens sampled uniformly at random, natural text can often be compressed, possibly allowing language models to copy longer strings even with limited memory. To test whether this matters, in 6(b) we conduct the same experiment as above but randomly shuffle the order of the words in the strings. We find that when we shuffle the words, both GSSMs and transformers perform worse on the task, but the effect is more stark for GSSMs. Even the largest GSSM now gets zero accuracy on strings of length 300. This suggests that when the input is more difficult to compress, the GSSM suffers due to its fixed size state.
@@ 4.3 Retrieval from the input context
We first consider a “phone-book” experiment where we provide a synthetic phone-book to the model and ask it to return the phone number when given a name. We generate the phone-book by randomly sampling $L$ names and their associated phone number. One line of this phone-book looks like “John Powell: 609-323-$7$”. Our prompt to the model consists of the phone-book, two few-shot examples and a question asking for the phone number of a randomly sampled name from the phone-book. 0(c) reports the accuracy obtained by the pretrained transformers and GSSMs while varying the size of the phone-book $L.$ We observe that even the smallest transformer (410M parameters) outperforms the largest GSSMs (2.8B parameters) when the phone-book size is long enough ($L\geq 70$). This shows that in retrieval tasks which require access to the whole context, GSSMs struggle to store the relevant information in their fixed-size state.
@@
In this experiment, we compare the 2.8B parameter Mamba and transformer models [^4] In our experiments, smaller models were unable to achieve reasonable and consistent performance on this dataset., on the SQuAD question-answering dataset (Rajpurkar et al., 2018). This dataset provides text paragraphs together with a few questions regarding the text. We probe the models to answer the question by providing a single demonstration of a question/answer pair (corresponding to the same text) before giving the target question. We bin the paragraphs according to their lengths, and report the F1 score as a function of the paragraph length for both models in 6(c). We observe that while for short paragraphs, both the Pythia transformer and Mamba achieve comparable performance, the performance of Mamba degrades more quickly with the paragraph length, while the transformer-based model maintains a similar accuracy even for longer texts. This result shows that the fixed-memory of GSSMs also limits their performance on standard natural tasks.
@@ 5 Related Work
Other papers have previously demonstrated the capacity of transformers to leverage the entire input context for tasks like retrieval, question answering, and in-context learning (Devlin et al., 2018; Raffel et al., 2020; Petroni et al., 2020; Brown et al., 2020; Liu et al., 2023b). Another line of work has studied the “induction head” mechanism in transformers that performs a retrieval operation much like the one we observe for copying (Olsson et al., 2022). But, to our knowledge, there is not a comparison in related work between transformers and GSSMs of similar quality on these tasks.
@@ 6 Discussion
We have demonstrated through theory and experiments that transformers are better than GSSMs at copying from their input context. However, we emphasize that state space models have many advantages over transformers. The memory and computational complexity of GSSMs does not increase with the input length, which is ideal for training and inference on long inputs. Additionally, state space models such as RNNs are better at tracking state variables across long sequences (Liu et al., 2023a), which may be useful for generating long consistent text. Importantly, language processing in the human brain appears to be much more similar to how state space models process language (Tikochinski et al., 2024). We therefore believe that future work should focus on building hybrid architectures that endow state space models with an attention-like mechanism, allowing them to retrieve relevant pieces of text from their input. Indeed, humans have an incredibly limited capacity for memorizing sequences (Miller, 1956), but can translate entire novels if we allow them to look back at the text (Shelton, 1612).
@@ A.1 Positional encodings in the transformers
  • the NoPE scheme (Kazemnejad et al., 2023) where no positional information is added to any of the attention scores (7(a)). This architecture choice helps to get better length generalization in multiple tasks including the copy task.
  • the ALiBi scheme (Press et al., 2021) which biases the attention scores with a penalty that is proportional to their distance (7(b)). $m$ is a head-specific slope fixed before training.
  • the Hard-ALiBi scheme introduced in Section 2 which has $M$ masked attention heads where we explicitly force the model to attend to their directly previous tokens and $H-M$ heads set to be NoPE attention heads. In 7(c), we display the case where we have $M=4$ masked heads: in the first head, the tokens just attend to themselves; in the second head, the tokens attend to themselves and to previous ones; in the third head, the tokens attend to themselves, the previous ones and the second preceding tokens. The remaining $H-M$ heads are set to NoPE.
@@ A.2 Pretraining and evaluation details
In Section 3, at each epoch, we sample online a batch size of size 64. We fill the context with examples so we choose a context length ($C=420$ for all the experiments except 0(a) where we set $C=620$) and pack as many examples as possible to fit this context. So in our case, one sample contains many instances. We run the experiments for 15 epochs for both transformers and Mamba while for LSTMs we need 300 epochs. All methods are trained with the AdamW optimizer (Loshchilov & Hutter, 2017) with learning rate 5e-5, a linear rate decay schedule, 300 steps of warmup and default weight decay of 1e-1. Finally, to train all the models, we use the next-token prediction loss but we apply a mask on the input instance so that we only penalize the model whenever it makes a mistake on the labels (and not on the inputs and labels jointly).
@@ B.1 Data efficiency on the copy task
Figure 9: (a) Copying long strings: character-level accuracy. Here we train models to copy strings of length $\leq 300$ and evaluate character-level accuracy on strings of length 300. Transformers train much faster than GSSMs. Mamba has a more progressive learning curve than in 0(a). An LSTM cannot even learn the task within this number of samples at the character level. (b) Copying short strings: string-level accuracy. Here we train models to copy strings of length $\leq 30$ and evaluate character-level accuracy on strings of length 30. Transformers train much faster than GSSMs. Compared to 0(a), we see that Mamba needs way less samples in order to learn to copy length-30 strings. An LSTM can learn to copy but requires 100x more training examples. (c) Copying short strings: character-level accuracy. Here we train models to copy strings of length $\leq 30$ and evaluate character-level accuracy on strings of length 30 and report the character-level accuracy.
In this section, we provide additional plots to complement the data efficiency experiment from 0(a). We want to highlight the following points:
  • in 0(a), we see a sharp transition for the Mamba learning curve. However, 8(a) shows that the learning process is more smooth at the character level. Besides, LSTMs are not able to learn the copy on length-300 strings even at the character level.
  • We consider the experiment of learning to copy much shorter strings namely strings with length $\leq 30$. 8(b) shows that the gap in terms of training examples between transformers and Mamba is much smaller i.e. Mamba only needs 10x more data. Besides, we see that the LSTM is able to learn the copy task but it needs 100x more data than transformers.
@@ B.2 Pre-trained models on the uniform copy task
This setting is a more extreme version of 6(b) since the strings are more random: in 6(b), the order of the nouns were random but the nouns were English nouns while in 6(b), the strings are totally random. In Figure 10, we see a clear separation between the transformers and Mamba models with the smallest Pythia outperforming the largest Mamba. However, compared to 6(b), the Pythia performance is much higher since the 1.4B model able to get almost 100% accuracy.
@@ C.1 Technical Lemmas
Proof: Using the previous claim, togehter with Lemma C.3, we get that:
@@ Appendix D Proofs - Lower Bound
Let $x_{n^{\prime}+1},\dots,x_{n^{\prime}+n}$ be the outputs of $H_{u,r}$. We need to show that there exist functions $G_{1},\dots,G_{n}$ s.t. $H_{u,r}(x_{1},\dots,x_{n^{\prime}})=G(S_{n^{\prime}}(x_{1},\dots,x_{n}))$. We give the following recurive definition:

### v2

@@ 2.2 Transformers can copy inputs of exponential length
In this section, we show that transformers can implement the copy operation for input sequences with length exponential in the number of heads. Namely, we construct a transformer with two blocks that gets small error on the copy task.
@@ 3.1 Experimental setup
We now provide a brief overview of our experimental setup. Further details may be found in Appendix A. Code and data available at: https://github.com/sjelassi/transformers_ssm_copy
@@ 3.2 Data efficiency on the copy task
Note that the sharp changes in accuracy displayed in 1(a) are due to the log-scaled x-axis and choice of string-level accuracy as a metric. In 9(a), we report the character-level accuracy, which yields smoother curves demonstrating the learning process of GSSMs. Regarding LSTMs, we find that they do not manage to learn on length-300 strings even at the character level. In 9(b), we show that LSTMs are able to learn to copy on shorter strings and that string length is the bottleneck.
@@ 3.3 Length generalization on the copy task
Here, we train all models on sequences of $\leq 50$ tokens, and test them on sequences of up to $1000$ tokens, reporting string-level accuracy. As seen in 1(b), all models are able to (eventually) solve the task in-distribution on lengths of $\leq 50$, but transformer-based models display much better generalization to longer inputs compared to GSSMs. Namely, we observe that the performance of the GSSMs (LSTM and MAMBA) drops to zero almost immediately when increasing the input length, while the performance of transformers decays much more gradually with length.
@@
When looking at the relative performance of different transformer models in 1(b), it becomes clear that the positional encoding is important to length generalization. Specifically, the ALiBi and NoPE transformers dramatically outperform the RoPE model on longer inputs. This is likely because the sinusoidal embeddings of RoPE create a more dramatic change than the decay of ALiBi or NoPE when we go to longer inputs.
@@
To test our understanding of how transformers learn to copy, we now consider swapping in the Hard-ALiBi positional encoding that we used in our theoretical construction of hash-based copying (introduces in Subsection 2.2 and illustrated in Figure 8 in the Appendix). 1(b) shows that a transformer trained with Hard-ALiBi embedding on sequences of length $\leq 50$ achieves almost perfect length generalization up to sequences of length 1000. Note that this is well beyond the context length ever encountered in training.
@@ 3.4 Transformers learn to use n-gram hashing
Next, we attempt to determine whether the transformer trained on the copy task indeed applies the mechanism of storage and retrieval of n-grams. To do this, we evaluate the performance of a transformer with Hard-ALiBi positional encoding trained on the copy task when tested on a distribution of examples that intentionally contains duplicate n-grams. That is, we draw uniform sequences of tokens, and then randomly replace some n-gram with another n-gram that already appears in the sequence, such that each example always contains two copies of the same n-gram (typically followed by a different token). We use the Hard-Alibi model here since it performs the best for the copy task as showed in 1(a). Figure 4 shows the performance of the transformer for different choices of $n$. We observe that the transformer maintains roughly the same accuracy for $n\leq 4$, but that its accuracy starts dropping when the inputs contains duplicate sequences of 5 or more tokens. This suggests that the transformer relies on something like 5-gram retrieval to do the copy task. Figure 11 further strengthens this point. We report the performance of perfect n-gram models in the copy task and observe that the performance of Transformers enhanced with Hard-ALiBi matches with the one of a 5-gram model.
@@ 4.1 Setup
We compare these models by measuring their performance while varying the input instance length and consider two types of tasks: copy-based and information retrieval tasks. The copy-based tasks consist of presenting a random text to the model and asking it to copy the text. In the information retrieval tasks, we provide a text to the model and ask it a related question. These retrieval tasks can be seen as “selective copy”, since the model needs to copy a small chunk of the input text in order to respond to the question. To measure performance, we use the string-level accuracy in all the experiments except in 7(c) where we consider question answering and thus report the F1 score. We evaluate the models over 10 batches of size 64 for all the tasks except for question answering where we evaluate over 50 questions because the number of questions with a given context length is limited. Further details are in Appendix A.
@@ 4.2 Copying the input context
We first observe that pre-trained transformers outperform pre-trained GSSMs at copying long natural language strings. In 7(a), we randomly sample strings from the C4 dataset (Raffel et al., 2020) with varying number of tokens. Our prompt consists of two copies of the sampled string plus the first word of the string and we expect the model to complete the third copy. Even the smallest transformer model dramatically outperforms the largest GSSM. This happens even though the large GSSMs have enough bits in the state variable to potentially store the context. This confirms the idea that this is an architectual bias of transformers that makes it easier for them to copy from the context.
Unlike strings of tokens sampled uniformly at random, natural text can often be compressed, possibly allowing language models to copy longer strings even with limited memory. To test whether this matters, in 7(b) we conduct the same experiment as above but randomly shuffle the order of the words in the strings. We find that when we shuffle the words, both GSSMs and transformers perform worse on the task, but the effect is more stark for GSSMs. Even the largest GSSM now gets zero accuracy on strings of length 300. This suggests that when the input is more difficult to compress, the GSSM suffers due to its fixed size state.
@@ 4.3 Retrieval from the input context
We first consider a “phone-book” experiment where we provide a synthetic phone-book to the model and ask it to return the phone number when given a name. We generate the phone-book by randomly sampling $L$ names and their associated phone number. One line of this phone-book looks like “John Powell: 609-323-$7$”. Our prompt to the model consists of the phone-book, two few-shot examples and a question asking for the phone number of a randomly sampled name from the phone-book. 1(c) reports the accuracy obtained by the pretrained transformers and GSSMs while varying the size of the phone-book $L.$ We observe that even the smallest transformer (410M parameters) outperforms the largest GSSMs (2.8B parameters) when the phone-book size is long enough ($L\geq 70$). This shows that in retrieval tasks which require access to the whole context, GSSMs struggle to store the relevant information in their fixed-size state.
@@
In this experiment, we compare the 2.8B parameter Mamba and transformer models [^4] In our experiments, smaller models were unable to achieve reasonable and consistent performance on this dataset., on the SQuAD question-answering dataset (Rajpurkar et al., 2018). This dataset provides text paragraphs together with a few questions regarding the text. We probe the models to answer the question by providing a single demonstration of a question/answer pair (corresponding to the same text) before giving the target question. We bin the paragraphs according to their lengths, and report the F1 score as a function of the paragraph length for both models in 7(c). We observe that while for short paragraphs, both the Pythia transformer and Mamba achieve comparable performance, the performance of Mamba degrades more quickly with the paragraph length, while the transformer-based model maintains a similar accuracy even for longer texts. This result shows that the fixed-memory of GSSMs also limits their performance on standard natural tasks.
@@ 5 Related Work
Other papers have previously demonstrated the capacity of transformers to leverage the entire input context for tasks like retrieval, question answering, and in-context learning (Devlin et al., 2018; Raffel et al., 2020; Petroni et al., 2020; Brown et al., 2020; Liu et al., 2023b; Kamradt, 2023). Another line of work has studied the “induction head” mechanism in transformers that performs a retrieval operation much like the one we observe for copying (Olsson et al., 2022). But, to our knowledge, there is not a comparison in related work between transformers and GSSMs of similar quality on these tasks.
@@
Concurrently to our work, Akyürek et al. (2024); Grazzi et al. (2024); Park et al. (2024) studied the difference between Transformers and Mamba at in-context learning, which can be seen as a form of copying. In particular, Akyürek et al. (2024) finds that Transformers have an advantage over other architectures at this task because they have “n-gram heads”. Similarly to these works, we hint the limitations of SSMs in memory-intensive tasks such as copying because of their limited state size. We also show that Transformers can perform copying using the Hard-ALiBi positional encoding, which improves the model’s ability to learn $n$-gram matching.
@@ 6 Discussion
We have demonstrated through theory and experiments that transformers are better than GSSMs at copying from their input context. However, we emphasize that state space models have many advantages over transformers. The memory and computational complexity of GSSMs does not increase with the input length, which is ideal for training and inference on long inputs. Additionally, state space models such as RNNs are better at tracking state variables across long sequences (Liu et al., 2023a), which may be useful for generating long consistent text. Importantly, language processing in the human brain appears to be much more similar to how state space models process language (Tikochinski et al., 2024).
We therefore believe that future work should focus on building hybrid architectures that endow state space models with an attention-like mechanism, allowing them to retrieve relevant pieces of text from their input. Indeed, humans have an incredibly limited capacity for memorizing sequences (Miller, 1956), but can translate entire novels if we allow them to look back at the text (Shelton, 1612).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1 Positional encodings in the transformers
  • the NoPE scheme (Kazemnejad et al., 2023) where no positional information is added to any of the attention scores (8(a)). This architecture choice helps to get better length generalization in multiple tasks including the copy task.
  • the ALiBi scheme (Press et al., 2021) which biases the attention scores with a penalty that is proportional to their distance (8(b)). $m$ is a head-specific slope fixed before training.
  • the Hard-ALiBi scheme introduced in Section 2 which has $M$ masked attention heads where we explicitly force the model to attend to their directly previous tokens and $H-M$ heads set to be NoPE attention heads. In 8(c), we display the case where we have $M=4$ masked heads: in the first head, the tokens just attend to themselves; in the second head, the tokens attend to themselves and to previous ones; in the third head, the tokens attend to themselves, the previous ones and the second preceding tokens. The remaining $H-M$ heads are set to NoPE.
@@ A.2 Pretraining and evaluation details
In Section 3, at each epoch, we sample online a batch size of size 64. We fill the context with examples so we choose a context length ($C=420$ for all the experiments except 1(a) where we set $C=620$) and pack as many examples as possible to fit this context. So in our case, one sample contains many instances. We run the experiments for 15 epochs for both transformers and Mamba while for LSTMs we need 300 epochs. All methods are trained with the AdamW optimizer (Loshchilov & Hutter, 2017) with learning rate 5e-5, a linear rate decay schedule, 300 steps of warmup and default weight decay of 1e-1. For LSTMs and Mamba, we did a sweep over learning rates $\{5\mathrm{e}-5,1\mathrm{e}-4,5\mathrm{e}-4\}.$ Finally, to train all the models, we use the next-token prediction loss but we apply a mask on the input instance so that we only penalize the model whenever it makes a mistake on the labels (and not on the inputs and labels jointly).
@@ B.1 Data efficiency on the copy task
Figure 9: (a) Copying long strings: character-level accuracy. Here we train models to copy strings of length $\leq 300$ and evaluate character-level accuracy on strings of length 300. Transformers train much faster than GSSMs. Mamba has a more progressive learning curve than in 1(a). An LSTM cannot even learn the task within this number of samples at the character level. (b) Copying short strings: string-level accuracy. Here we train models to copy strings of length $\leq 30$ and evaluate character-level accuracy on strings of length 30. Transformers train much faster than GSSMs. Compared to 1(a), we see that Mamba needs way less samples in order to learn to copy length-30 strings. An LSTM can learn to copy but requires 100x more training examples. (c) Copying short strings: character-level accuracy. Here we train models to copy strings of length $\leq 30$ and evaluate character-level accuracy on strings of length 30 and report the character-level accuracy.
In this section, we provide additional plots to complement the data efficiency experiment from 1(a). We want to highlight the following points:
  • in 1(a), we see a sharp transition for the Mamba learning curve. However, 9(a) shows that the learning process is more smooth at the character level. Besides, LSTMs are not able to learn the copy on length-300 strings even at the character level.
  • We consider the experiment of learning to copy much shorter strings namely strings with length $\leq 30$. 9(b) shows that the gap in terms of training examples between transformers and Mamba is much smaller i.e. Mamba only needs 10x more data. Besides, we see that the LSTM is able to learn the copy task but it needs 100x more data than transformers.
@@ B.2 Pre-trained models on the uniform copy task
This setting is a more extreme version of 7(b) since the strings are more random: in 7(b), the order of the nouns were random but the nouns were English nouns while in 7(b), the strings are totally random. In Figure 10, we see a clear separation between the transformers and Mamba models with the smallest Pythia outperforming the largest Mamba. However, compared to 7(b), the Pythia performance is much higher since the 1.4B model able to get almost 100% accuracy.
B.3 Performance $n$-gram models at copying
n-gram length:  2    3    4    5
Transformers:  Hard-ALiBi
Figure 11: String-level copying accuracy obtained by perfect $n$-gram models and Transformers with Hard-ALiBi. Transformers performance matches the one of 5-gram model.
In Figure 11, we display the performance of perfect $n$-gram models in the copy task. To obtain these curves, we uniformly sample 128 strings over 3 seeds and report the probability there is a $n$-gram. This probability corresponds to the performance of a perfect $n$-gram model. We observe that Transformers enhanced with the Hard-ALiBi positional encoding have a performance close to a perfect $5$-gram model.
@@ C.1 Technical Lemmas
Proof: Using the previous claim, together with Lemma C.3, we get that:
@@ Appendix D Proofs - Lower Bound
Let $x_{n^{\prime}+1},\dots,x_{n^{\prime}+n}$ be the outputs of $H_{u,r}$. We need to show that there exist functions $G_{1},\dots,G_{n}$ s.t. $H_{u,r}(x_{1},\dots,x_{n^{\prime}})=G(S_{n^{\prime}}(x_{1},\dots,x_{n}))$. We give the following recursive definition: